the equation y=5a-3b+c, if a=12, b=4,and c=8 my answer was 40 is this correct?
Answer by checkley71(8403)

Y=5A-3B+C OR Y=5(12)-3(4)+8 OR Y=60-12+8 OR Y=68-12 OR 56 NOT 40